# Saving SWE-Bench: A Benchmark Mutation Approach for Realistic Agent Evaluation

### v3

@@ Abstract.
Current benchmarks for evaluating software engineering agents, such as SWE-Bench Verified, are predominantly derived from GitHub issues and fail to accurately reflect how developers interact with chat-based coding assistants in integrated development environments (IDEs). We posit that this mismatch leads to a systematic overestimation of agent’s capabilities in real-world scenarios, especially bug fixing. We introduce a novel benchmarking framework that transforms existing formal benchmarks into realistic user queries through systematic analysis of developer interaction patterns with chat-based agents. Our methodology is flexible and can be easily extended to existing benchmarks. In this paper, we apply our testing framework to SWE-Bench Verified, the TypeScript subset of Multi-SWE-Bench and a private benchmark, SWE-Bench C# and transform formal GitHub issue descriptions into realistic user-style queries based on telemetry analysis of a popular chat-based agent interactions. Our findings reveal that existing benchmarks significantly overestimate agent capabilities for some models by >>50% over baseline performance for public benchmarks and $\sim$10-16% for our internal benchmark. This work establishes a new paradigm for evaluating interactive chat-based software engineering agents through benchmark mutation techniques. Our code is available at: https://anonymous.4open.science/r/SWE-Bench-Mutated-ED8F
@@ 1. Introduction
4. Code Artifacts: We incorporate our approach into the official SWE-Bench GitHub repo and share our working implementation consisting of all the prompts we used and mutations generated for the SWE-Bench Verified dataset [^1] Implementation of SWE-Bench-Mutated available at: https://anonymous.4open.science/r/SWE-Bench-Mutated-ED8F. We hope that this allows the community to easily replicate our results and adopt this technique in future agent evaluations.
@@ 10. User Interface & UX Design
Figure 2. The prompt we use for labeling a user query into one of the 10 high-level categories we found in our data.
@@
    Apply Categorization to New Queries: In the second stage, we applied the categorization prompt shown in Figure 2 to assign each query to its most appropriate category. Examples of such queries are shown in Figure  4.
@@
Figure 3 shows the distribution of the resulting categories. We can see that bug fixing and error resolution is among the largest category of user interactions, accounting for $\sim$14% of all queries.
@@ 3.1.2. Understanding Bug-Fix User Queries
To further demonstrate the differences between the communication styles of users, we measure how often benchmarks and user queries from telemetry contain various key elements that make a bug report useful, such as file-path(s), reproduction code, test case(s), expected vs actual behavior and environment / configuration details, using prompt shown in Figure 7 with an LLM. Figure 6 shows a comparison of how often these elements show up in the queries contained within benchmarks, containing primarily GitHub issues, and user queries. We can see that users tend to communicate more with the error stack and file-paths compared to any other element in agent queries. In stark contrast, GitHub issues contain other elements like reproduction code, expected behavior, environment details as well. These features are largely missing from user queries. Our analysis reflects chat users’ preference for sharing more targeted information like file-paths, error stacks, rather than providing exhaustive details upfront as typically seen in GitHub issues.
@@
This analysis yielded 11 distinct communication templates that capture how developers typically report bugs to chat-based agents:
@@ 3.2. Mutation Methodology
Building on our template analysis, we developed a systematic methodology for transforming formal benchmark problems into queries matching real-world usage patterns. Figure 9 shows the high-level pipeline of our approach. We describe our mutation algorithm in more detail below and then give an example of a mutation generated by the algorithm for SWE-Bench.
@@ 4.1. Benchmarks
  • Multi-SWE-Bench (TypeScript): We mutate the TypeScript subset of Multi-SWE-Bench. We chose to focus only on the TypeScript subset (224 examples) due to the size of the overall dataset. We call the mutated version of this benchmark Multi-SWE-Bench (TypeScript)-Mutated.
For each of the above benchmarks, we apply the mutation prompt in Figure 11 and generate multiple possible mutations for each problem statement. For each example, we randomly pick one mutation from the set of possible mutations generated by the model.
@@ 5.1.1. SWE-Bench Verified Results
These results demonstrate that existing benchmarks significantly overestimate agent’s bug fixing capabilities. The consistent performance drops across models suggest that the gap is not simply due to model-specific limitations, but rather reflects fundamental differences in problem presentation and agent adaptation to realistic scenarios. We see a higher drop for GPT-4.1 compared to other models. This could be because this model perhaps requires more context, such as in a GitHub issue, to solve a given problem or is possibly over-fitted to SWE-Bench.
@@ 5.2.3. Multi-SWE-Bench (TypeScript) Results
Similar to SWE-Bench C#, Sonnet 4 also consumes a lot more tokens and steps for TypeScript bug-fixing tasks, unlike tasks in SWE-Bench Verified. GPT-4.1 also consumes a lot more steps and tokens for TypeScript compared to the other two benchmarks.
@@ 6. Limitations
While our study provides valuable insights into mutating benchmarks for realistic agent evaluation, several limitations should be acknowledged and present opportunities for future research.
@@
Unchanged Evaluation Methodology & Harness: Although we modify problem statements to reflect realistic user queries, the evaluation harness we use remains unchanged and rigid. Ideally, we would develop a better set of metrics that capture the nuances of human-agent interaction and better approximate user satisfaction with agent responses.
Despite these limitations, our work establishes a foundational approach toward realistic evaluation for bug-fixing capabilities in chat-based coding agents and provides an extensible framework that can be adapted to existing and new benchmarks.

### v4

@@ Abstract.
Current benchmarks for evaluating software engineering agents, such as SWE-Bench Verified, are predominantly derived from GitHub issues and fail to accurately reflect how developers interact with chat-based coding assistants in integrated development environments (IDEs). We posit that this mismatch leads to a systematic overestimation of agent’s capabilities in real-world scenarios, especially bug fixing. We introduce a novel benchmarking framework that transforms existing formal benchmarks into realistic user queries through systematic analysis of developer interaction patterns with chat-based agents. Our methodology is flexible and can be easily extended to existing benchmarks. In this paper, we apply our testing framework to SWE-Bench Verified, the TypeScript subset of Multi-SWE-Bench and a private benchmark, SWE-Bench C# and transform formal GitHub issue descriptions into realistic user-style queries based on telemetry analysis of a popular chat-based agent interactions. Our findings reveal that existing benchmarks significantly overestimate agent capabilities for some models by >>50% over baseline performance for public benchmarks and $\sim$10-16% for our internal benchmark. This work establishes a new paradigm for evaluating interactive chat-based software engineering agents through benchmark mutation techniques. Our code is available at: https://github.com/microsoft/SWE-Bench-Mutated-CAIN26
conference: 5th International Conference on AI Engineering – Software Engineering for AI; April 12–13, 2026; Rio de Janeiro, Brazil
@@ 1. Introduction
4. Code Artifacts: We incorporate our approach into the official SWE-Bench GitHub repo and share our working implementation consisting of all the prompts we used and mutations generated for the SWE-Bench Verified dataset [^1] Implementation of SWE-Bench-Mutated available at: https://github.com/microsoft/SWE-Bench-Mutated-CAIN26. We hope that this allows the community to easily replicate our results and adopt this technique in future agent evaluations.
@@ 10. User Interface & UX Design
Figure 2. The prompt we use for labeling a user query into one of the 10 high-level categories in our data.
@@
    Apply Categorization to New Queries: In the second stage, we applied the categorization prompt shown in Figure 2 to assign each query to its most appropriate category. Figure  4 shows examples of such queries.
@@
Figure 3 shows the distribution of the resulting categories. We can see that bug fixing and error resolution is among the largest categories, accounting for $\sim$14% of all queries.
@@ 3.1.2. Understanding Bug-Fix User Queries
To further demonstrate the differences between the communication styles of users, we measure how often benchmarks and user queries from telemetry contain various key elements that make a bug report useful, such as file-path(s), reproduction code, test case(s), expected vs actual behavior and environment / configuration details, using prompt shown in Figure 7 with an LLM. Figure 6 shows a comparison of how often these elements show up in the queries contained within benchmarks, containing primarily GitHub issues, and user queries. We can see that users tend to communicate more with the error stack and file-paths compared to other elements in agent queries. In stark contrast, GitHub issues contain reproduction code, expected behavior, environment details, etc. as well. These features are largely missing from user queries. Our analysis reflects chat users’ preference for sharing more targeted information like file-paths, error stacks, rather than providing exhaustive details upfront as typically seen in GitHub issues.
@@
This yielded 11 distinct communication templates that capture how developers report bugs to chat-based agents:
@@ 3.2. Mutation Methodology
Building on our template analysis, we developed a systematic methodology for transforming formal benchmark problems into queries matching real-world usage patterns. Figure 9 shows the high-level pipeline of our approach. We describe our mutation algorithm in more detail below and then give an example of a mutation generated for SWE-Bench.
@@ 4.1. Benchmarks
  • Multi-SWE-Bench (TypeScript): We mutate the TypeScript subset of Multi-SWE-Bench. We chose to focus only on a subset (224 examples) due to the size of the overall dataset. We call the mutated version of this benchmark Multi-SWE-Bench (TypeScript)-Mutated.
For each benchmarks, we apply the mutation prompt in Figure 11 and generate multiple mutations for each problem statement. For each example, we randomly pick one mutation from the set of possible mutations generated by the model.
@@ 5.1.1. SWE-Bench Verified Results
These results demonstrate that existing benchmarks significantly overestimate agent’s bug fixing capabilities. The consistent performance drops across models suggest that the gap is not simply due to model-specific limitations, but rather reflects fundamental differences in problem presentation and agent adaptation to realistic scenarios. We see a higher drop for GPT-4.1 compared to other models. This could be because this model perhaps requires more context, such as in a GitHub issue, to solve a given problem or is possibly over-fitted.
@@ 5.2.3. Multi-SWE-Bench (TypeScript) Results
Similar to SWE-Bench C#, Sonnet 4 also consumes a lot more tokens and steps for TypeScript bug-fixing tasks, unlike tasks in SWE-Bench Verified. GPT-4.1 also takes more steps and tokens compared to other benchmarks.
@@ 6. Limitations
While our study provides valuable insights into mutating benchmarks for realistic agent eval, we acknowledge several limitations and present opportunities for future research.
@@
Unchanged Evaluation Methodology & Harness: Though we modify problem statements to reflect realistic user queries, the eval harness we use remains unchanged and rigid. Ideally, we would develop a better set of metrics that capture the nuances of human-agent interaction and better approximate user satisfaction with agent responses.
Despite these limitations, our work establishes a foundational approach toward realistic eval for bug-fixing capabilities in chat-based coding agents and provides an extensible framework that can be adapted to existing and new benchmarks.